brenda purchased a business for 120,000. and received s building fmv 80,000. land fmv 10,000. and furniture and fixtures fmv 30,000. she spent 10,000. on legal fees to make thr purchase what is the basis for these items
Answer #1

Total acquisition cost = 120,000 + 10,000 = 130,000

Total fmv = 80,000+10,000+30,000 = 120,000

Building (80,000/120,000)*130,000 86,667
Land (10,000/120,000)*130,000 10,833
Furniture and fixtures (30,000/120,000)*130,000 32,500